### v1

# On shallow feedforward neural networks with inputs from a topological space

Abstract. We study feedforward neural networks with inputs from a topological space (TFNNs). We prove a universal approximation theorem for shallow TFNNs, which demonstrates their capacity to approximate any continuous function defined on this topological space. As an application, we obtain an approximative version of Kolmogorov’s superposition theorem for compact metric spaces.
Mathematics Subject Classifications: 41A30, 41A65, 68T05
Keywords: feedforward neural network, universal approximation theorem, density, topological vector space, Tauber-Wiener function, Kolmogorov’s superposition theorem
1. Introduction
Neural networks are fundamental to contemporary machine learning and artificial intelligence, providing robust methods for tackling intricate challenges. Among the different neural network designs, the multilayer feedforward perceptron (MLP) is particularly prominent and essential. The MLP is valued for its capability to model complex, nonlinear functions and execute various tasks, including classification, regression, and pattern recognition.
This architecture consists of a limited number of sequential layers: an input layer at the beginning, an output layer at the end, and several hidden layers in between. Information progresses from the input layer through the hidden layers to the output layer. In this framework, each neuron in a layer receives inputs from the previous layer, applies specific weights, adds a bias, and then processes the result through an activation function. This activation function introduces non-linearity, allowing the model to learn and capture intricate patterns. The output from one layer’s neurons serves as the input for the neurons in the next layer, continuing this sequence until the final output is generated by the output layer.
The most basic form of an MLP features just one hidden layer. In this setup, each output neuron calculates a function expressed as
$\sum_{i=1}^{r}c_{i}\sigma(\mathbf{w}^{i}\cdot\mathbf{x}-\theta_{i}),$ (1.1)1.1( 1.1 )
where $\mathbf{x}=(x_{1},...,x_{d})$ represents the input vector, $r$ is the number of neurons in the hidden layer, $\mathbf{w}^{i}$ are weight vectors in $\mathbb{R}^{d}$, $\theta_{i}$ are thresholds, $c_{i}$ are coefficients, and $\sigma$ is the activation function, a real univariate function.
The theoretical underpinning of neural networks is rooted in the universal approximation property (UAP), sometimes referred to as the density property. This principle states that a neural network with a single hidden layer can approximate any continuous function over a compact domain to any desired level of precision. Specifically, the set $span\{\sigma(\mathbf{w\cdot x}-\theta):\theta\in\mathbb{R},\mathbf{w}\in\mathbb{R}^{d}\}$, which comprises functions defined in the format of equation (1.1), is dense in $C(K)$ for every compact set $K\subset\mathbb{R}^{d}$. Here $C(K)$ represents the space of real-valued continuous functions on $K$. This important result in neural network theory is known as the universal approximation theorem (UAT).
Extensive research has investigated the UAT across various activation functions $\sigma$, examining how different choices influence the approximation capabilities of neural networks. The most general result in this area was obtained by Leshno, Lin, Pinkus and Schocken [14]. They proved that a continuous activation function $\sigma$ possesses the UAP if and only if it is not a polynomial. This result demonstrates the effectiveness of the single hidden layer perceptron model across a wide range of activation functions $\sigma$. It should be noted that, the universal approximation theorem in [14] was shown to apply to a broader class of activation functions beyond just continuous ones, including activation functions that may have discontinuities on sets of Lebesgue measure zero. However, this paper will specifically concentrate on continuous activation functions. For a thorough, step-by-step proof of this theorem, refer to [19, 20].
In the past, it was commonly accepted and highlighted in numerous studies that attaining the universal approximation property necessitate large networks with a substantial number of hidden neurons (see, e.g., [4, Chapter 6.4.1]). In the above-mentioned earlier works, the number of hidden neurons was regarded as unbounded. However, more recent research [5, 6, 7] has demonstrated that neural networks using certain non-explicit but practically computable activation functions can approximate any continuous function over any compact set to any desired level of accuracy, even with a minimal and fixed number of hidden neurons.
Note that the inner product $\mathbf{w}^{i}\cdot\mathbf{x}$ in (1.1) represents a linear continuous functional on $\mathbb{R}^{d}$. Conversely, by Riesz representation theorem, every linear functional on $\mathbb{R}^{d}$ is of the form $\mathbf{w\cdot x}$, where $\mathbf{w}\in\mathbb{R}^{d}$ and $\mathbf{x}=(x_{1},...,x_{d})$ is the variable (see [21, Theorem 13.32]). Linear continuous functionals constitute a significant subclass in $C(\mathbb{R}^{d})$, denoted here by $\mathcal{L}(\mathbb{R}^{d})$. Thus, UAT asserts that for certain activation functions $\sigma$ and any compact set $K\subset\mathbb{R}^{d}$, the set
$\mathcal{M}(\sigma)=span\{\sigma(f(x)-\theta):f\in\mathcal{L}(\mathbb{R}^{d}),\theta\in\mathbb{R}\}$
is dense in $C(K)$. This observation tells the following generalization of single hidden layer networks from $\mathbb{R}^{d}$ to any topological space $X,$ where $\mathcal{L}(\mathbb{R}^{d})$ is replaced with a fixed family of functions (which need not be linear) from $C(X)$. We refer to such a family as a basic family for the feedforward neural networks with inputs from a topological space (TFNNs). If $\mathcal{A}(X)\subset C(X)$ is a basic family, then the architecture of a single hidden layer TFNN can be described as follows:
  • Input Layer: This layer consists of an element $x\in X$, where $X$ is an arbitrary topological space.
[…]
  • Output layer: Each neuron in this layer receives weighted signals from each neuron in the hidden layer, sums them up, and produces the final output value.
This architecture significantly extends the traditional feedforward neural networks. When $X=\mathbb{R}^{d}$ and $\mathcal{A}(X)=\mathcal{L}(\mathbb{R}^{d})$, the input $x$ represents a $d$-dimensional vector. In this very special case, the layer contains $d$ traditional neurons, each receiving an input signal $x_{1},x_{2},...,x_{d}\in\mathbb{R}$, respectively. Note that in the aforementioned architecture, the element $x\in X$ carries all the information of the input layer. This structure enables the network to accommodate a wide diversity of input types. In general, a single hidden layer TFNN computes a function of the form
$\sum_{i=1}^{r}c_{i}\sigma(w_{i}f_{i}(x)-\theta_{i}),$ (1.2)1.2( 1.2 )
[…]
The aim of this paper is to show that for a broad a class activation functions $\sigma$, neural networks of the form presented in (1.2) can approximate any continuous function on a compact subset $K\subset X$ with arbitrary precision. In other words, the set
$\mathcal{N}(\sigma)=span\{\sigma(wf(x)-\theta):f\in\mathcal{A}(X);~{}w,\theta\in\mathbb{R}\}$
is dense in $C(K)$ for every compact set $K\subset X$. As an application of this result, we will derive an approximative version of the Kolmogorov superposition theorem (KST) for compact metric spaces, where outer (non-fixed and generally nonsmooth) functions are substituted with a fixed ultimately smooth function.
It should be noted that the UAP of neural networks operating between Banach spaces has been explored in various studies. For example, in [24], the fundamentality of ridge functions was established in a Banach space and subsequently applied to shallow networks with a sigmoidal activation function (see also [15]). In [2], the authors showed that any continuous nonlinear function mapping a compact set $V$ in a Banach space of continuous functions $C(K_{1})$ into $C(K_{2})$ can be approximated arbitrarily well by shallow feedforward neural networks. Here $K_{1}$ and $K_{2}$ represent two compact sets in an abstract Banach space $X$ and the Euclidean space $\mathbb{R}^{d}$, respectively. In [16], this approach was extended to deep neural networks and referred to as DeepONet. In [13], the authors examined DeepONet within the context of an encoder-decoder network framework, investigating its approximation properties when the input space is a Hilbert space. In [11], quantitative estimates (i.e., convergence rates) for the approximation of nonlinear operators using single-hidden layer networks in infinite-dimensional Banach spaces were provided, extending some previous results from the finite-dimensional case.
The UAP of infinite-dimensional neural networks, with inputs from Fréchet spaces and outputs from Banach spaces, was established in [1]. In [3], the scope of this architecture was extended by proving several universal approximation theorems for quasi-Polish input and output spaces.
In [12], universal approximation theorems were obtained for neural operators (NOs) and mixtures of neural operators (MoNOs) acting between Sobolev spaces. More precisely, it was shown that any non-linear continuous operator acting between Sobolev spaces $H^{s_{1}}$ and $H^{s_{2}}$ can be uniformly approximated over any compact set $K\subset$ $H^{s_{1}}$ with arbitrary accuracy $\varepsilon$ using NOs and MoNOs: $H^{s_{1}}\rightarrow H^{s_{2}}$. Moreover, the quantitative results of [12] estimate the depth, width, and rank of the neural operators in terms of the radius of $K$ and $\varepsilon$.
Recent research has demonstrated the universal approximation theorem (UAT) for various hypercomplex-valued neural networks, including complex-, quaternion-, tessarine-, and Clifford-valued networks, as well as more general vector-valued neural networks (V-nets) defined over a finite-dimensional algebra (see [25] and references therein). We hope that the results of this paper will stimulate further exploration of these neural networks, particularly with outputs from these and other general spaces.
2. Main results
In this section, we analyze the conditions under which shallow networks with inputs from a topological space possess the universal approximation property.
Assume $X$ is an arbitrary topological space. In the sequel, in $C(X)$, we will use the topology of uniform convergence on compact sets. This topology is induced by the seminorms
$\left\|g\right\|_{K}=\max_{x\in K}\left|g(x)\right|,$
where $K$ are compact sets in $X$. A subbasis at the origin for this topology is given by the sets
$U(K,r)=\left\{g\in C(X):\left\|g\right\|_{K}<r\right\},$
where $K\subset X$ is compact and $r>0$. A sequence (or net) $\{g_{n}\}$ in this topology converges to $g$ iff $\left\|g_{n}-g\right\|_{K}\rightarrow 0$ for every compact set $K\subset X$. Thus, in what follows, when we say that $B$ is dense in $C(X)$, we will mean that $B$ is dense with respect to the aforementioned topology of uniform convergence on compact sets.
We say that a subclass $\mathcal{A}(X)\subset C(X)$ holds the $D$-property if the set
$S=span\left\{u\circ v:u\in C(\mathbb{R}),v\in\mathcal{A}(X)\right\}$ (2.1)2.1( 2.1 )
is dense in $C(X)$.
In what follows, we will use activation functions $\sigma:\mathbb{R}\rightarrow\mathbb{R}$ (whether continuous or discontinuous) with the property that the $span\{\sigma(wx-\theta):w\in\mathbb{R},\theta\in\mathbb{R}\}$ is dense in every $C[a,b]$. Such functions are called Tauber-Wiener (TW) functions (see [2]).
Theorem 2.1. Assume $X$ is a topological space, $\mathcal{A}(X)$ is a subclass of $C(X)$ with the $D$-property and $\sigma:\mathbb{R}\rightarrow\mathbb{R}$ is a TW function. Then for any $\varepsilon>0$, any compact set $K\subset X$ and any function $g\in C(K)$ there exist $r\in\mathbb{N}$, $f_{i}\in\mathcal{A}(X)$, $c_{i},w_{i},\theta_{i},\in\mathbb{R}$, $i=1,...,r$, such that
$\max_{x\in K}\left|g(x)-\sum_{i=1}^{r}c_{i}\sigma(w_{i}f_{i}(x)-\theta_{i})\right|<\varepsilon.$
That is, TFNNs with inputs from $X$ is dense in $C(X)$.
Proof. Take any $\varepsilon>0$, any compact set $K\subset X$ and any function $g\in C(K)$. Since $\mathcal{A}(X)$ has the $D$-property, there exist finitely many functions $u_{i}\in C(\mathbb{R})$ and $v_{i}\in C(X)$ such that
$\left|g(x)-\sum_{i=1}^{n}u_{i}(v_{i}(x))\right|<\varepsilon/2,$ (2.2)2.2( 2.2 )
for all $x\in K$.
Since $v_{i}$ are continuous, the images $v_{i}(K)$ are compact sets in $\mathbb{R}$. Set $V=\cup_{i=1}^{n}v_{i}(K)$. Note that $V$ is also compact.
Since $\sigma$ is a TW function, each continuous univariate function $u_{i}(t)$, $t$ $\in V$, can be approximated by single hidden layer networks with the activation function $\sigma$. Thus, there exist coefficients $c_{ij},w_{ij},\theta_{ij}\in\mathbb{R}$, $1\leq i\leq n$, $1\leq j\leq k_{i}$, such that
$\left|u_{i}(t)-\sum_{j=1}^{k_{i}}c_{ij}\sigma(w_{ij}t-\theta_{ij})\right|<\varepsilon/2n$
for all $t\in V$. Therefore,
$\left|u_{i}(v_{i}(x))-\sum_{j=1}^{k_{i}}c_{ij}\sigma(w_{ij}v_{i}(x)-\theta_{ij})\right|<\varepsilon/2n$ (2.3)2.3( 2.3 )
for each $i=1,...,n$, and all $x\in K$. It follows from (2.2) and (2.3) that
$\left|g(x)-\sum_{i=1}^{n}\sum_{j=1}^{k_{i}}c_{ij}\sigma(w_{ij}v_{i}(x)-\theta_{ij})\right|<\varepsilon$
for any $x\in K$. This completes the proof of Theorem 2.1.
Remark. Theorem 2.1 generalizes existing universal approximation theorems for traditional feedforward neural networks. This is because in traditional networks, the space of linear continuous functionals on $\mathbb{R}^{d}$ serves as the basic family $\mathcal{A}(X)$, which clearly satisfies the $D$-property.
Note that, in particular, $X$ may be a topological vector space. For such a space, $X^{\ast}$ denotes the continuous dual of $X$, which is the space of linear continuous functionals defined on $X$. The following theorem is based on Theorem 2.1.
Theorem 2.2. Assume $X$ is a locally convex topological vector space (in particular, a normed space) and $\sigma$ is a continuous univariate function that is not a polynomial. Then for any $\varepsilon>0$, any compact set $K\subset X$ and any function $g\in C(K)$ there exist $r\in\mathbb{N}$, $f_{i}\in X^{\ast}$, $c_{i},\theta_{i}\in\mathbb{R}$, $i=1,...,r$, such that
$\max_{x\in K}\left|g(x)-\sum_{i=1}^{r}c_{i}\sigma(f_{i}(x)-\theta_{i})\right|<\varepsilon.$
The proof of this theorem relies on Theorem 2.1 and the following two facts.
Fact 1. The space $X^{\ast}$ possesses the $D$-property.
Let us prove this fact. Specifically, this property holds if in (2.1), instead of all $u\in C(\mathbb{R})$, we take the single function $u(t)=e^{t}$. That is, we claim that the set
$S=span\{e^{r(x)}:r\in X^{\ast}\}.$
is dense in $C(K)$ for every compact set $K\subset X$.
Indeed, first it is not difficult to see that $S$ is a subalgebra of $C(X)$. To see this, note that for any $r_{1},r_{2}\in X^{\ast}$
$e^{r_{1}(x)}e^{r_{2}(x)}=e^{r_{1}(x)+r_{2}(x)}\in E\text{,}$
since $r_{1}+r_{2}\in X^{\ast}$. Therefore, the linear space $S$ is closed under multiplication, indicating that $S$ is an algebra.
Second, if $r$ is the zero functional, then $e^{r(x)}=1$, showing that $S$ contains all constant functions.
Now since $X$ is locally convex, the Hahn-Banach extension theorem holds. It is a consequence of this theorem that for any distinct points $x_{1},x_{2}\in X$ there exists a functional $r\in X^{\ast}$ such that $r(x_{1})\neq r(x_{2})$ (see [22, Theorem 3.6]). Hence, the algebra $S$ separates points in $X$. By the Stone-Weierstrass theorem [23], for any compact $K\subset X$, the algebra $S$ restricted to $K$ is dense in $C(K)$. In other words, the space $X^{\ast}$ has the $D$-property.
Fact 2. A continuous nonpolynomial activation function $\sigma$ is a TW function.
This fact follows from the main result of [14] that a continuous nonpolynomial activation function provides the universal approximation property for traditional single hidden layer networks.
Let us now we apply Theorem 2.1 to derive an approximative version of the renowned Kolmogorov superposition theorem (KST) for compact metric spaces. KST [10] states that for the unit cube $\mathbb{I}^{d},~{}\mathbb{I}=[0,1],~{}d\geq 2,$ there exist $12d+$ 1 functions $\{s_{q}\}_{q=1}^{2d+1}\subset C(\mathbb{I}^{d})$ of the form
$s_{q}(x_{1},...,x_{d})=\sum_{p=1}^{d}\varphi_{pq}(x_{p}),~{}\varphi_{pq}\in C(\mathbb{I}),~{}p=1,...,d,~{}q=1,...,2d+1,$
such that each function $f\in C(\mathbb{I}^{d})$ admits the representation
$f(\mathbf{x})=\sum_{q=1}^{2d+1}g_{q}(s_{q}(\mathbf{x})),~{}\mathbf{x}=(x_{1},...,x_{d})\in\mathbb{I}^{d},~{}g_{q}\in C({{\mathbb{R)}}}.$
This surprising and deep result, which solved (negatively) Hilbert’s 13-th problem, has been improved and generalized in several directions. For detailed information about KST, including its refinements, various variants, and generalizations, see the monographs [9, Chapter 1] and [7, Chapter 4]. The relevance of KST to neural networks, along with its theoretical and computational aspects, has been extensively discussed in the neural network literature (see, e.g., [8] and references therein).
Ostrand [18] extended KST to general compact metric spaces as follows.
Theorem 2.3. (Ostrand [18]). For $p=1,2,...,n$, let $X_{p}$ be a compact metric space of finite dimension $d_{p}$ and let $m=\sum_{p=1}^{n}d_{p}.$ There exist universal continuous functions $\psi_{pq}:X_{p}\rightarrow[0,1],$ $p=1,...,n,$ $q=1,...,2m+1,$ such that every continuous function $g$ defined on $\Pi_{p=1}^{n}X_{p}$ is representable in the form
$g(x_{1},...,x_{n})=\sum_{q=1}^{2m+1}h_{q}(\sum_{p=1}^{n}\psi_{pq}(x_{p})),$
where $h_{q}$ are continuous functions depending on $g$.
It follows from this theorem that for the metric space $X=\Pi_{p=1}^{n}X_{p}$ the family of $12m+$ 1 functions
$\mathcal{K}(X)=\left\{\sum_{p=1}^{n}\psi_{pq}(x_{p}):q=1,...,2m+1\right\}$
satisfies the $D$-property in $C(X)$.
Let now $\sigma$ be a specific infinitely differentiable TW function with the property that, for any interval $[a,b]$, the set $\Lambda=\{\sigma(wx-\theta):w,\theta\in\mathbb{R}\}$ is dense in $C[a,b]$. Note that this is not the linear span of the functions $\sigma(wx-\theta)$, but rather a very narrow subclass of it. Such functions $\sigma$ do indeed exist.
To show this, let $\alpha$ be any positive real number. Divide the interval $[\alpha,+\infty)$ into the segments $[\alpha,2\alpha],$ $[2\alpha,3\alpha],...$. Let $\{p_{n}(t)\}_{n=1}^{\infty}$ be the sequence of polynomials with rational coefficients defined on $[0,1].$ We construct $\sigma$ in two stages. In the first stage, we define $\sigma$ on the closed intervals $[(2m-1)\alpha,2m\alpha],$ $m=1,2,...$ as the function
$\sigma(t)=p_{m}(\frac{t}{\alpha}-2m+1),\text{ }t\in[(2m-1)\alpha,2m\alpha],$
or equivalently,
$\sigma(\alpha t+(2m-1)\alpha)=p_{m}(t),\text{ }t\in[0,1].$ (2.4)2.4( 2.4 )
In the second stage, we extend $\sigma$ to the intervals $(2m\alpha,(2m+1)\alpha),$ $m=1,2,...,$ and $(-\infty,\alpha)$, maintaining the $C^{\infty}$ property.
For any univariate function $h\in C[0,1]$ and any $\varepsilon>0$ there exists a polynomial $p(t)$ with rational coefficients such that
$\left|h(t)-p(t)\right|<\varepsilon,$
for all $t\in[0,1].$ This together with (2.4) mean that
$\left|h(t)-\sigma(\alpha t-s)\right|<\varepsilon,$ (2.5)2.5( 2.5 )
for some $s\in\mathbb{R}$ and all $t\in[0,1].$
Using linear transformation it is not difficult to go from $[0,1]$ to any finite closed interval $[a,b]$. Indeed, let $u\in C[a,b]$, $\sigma$ be constructed as above and $\varepsilon$ be an arbitrarily small positive number. The transformed function $h(t)=u(a+(b-a)t)$ is well defined on $[0,1]$ and we can apply the inequality (2.5). Now using the inverse transformation $t=\frac{x-a}{b-a}$, we can write that
$\left|u(x)-\sigma(wx-\theta)\right|<\varepsilon,$ (2.6)2.6( 2.6 )
for all $x\in[a,b]$, where $w=\frac{\alpha}{b-a}$ and $\theta=\frac{\alpha a}{b-a}+s$.
We define activation functions $\sigma$ as superactivation functions if they satisfy (2.6) for any $u\in C[a,b]$, $\varepsilon>0$, and some $w,\theta\in\mathbb{R}$. These functions demonstrate that shallow networks can approximate univariate continuous functions with the minimal number of hidden neurons; in fact, a single hidden neuron is sufficient. Similar activation functions $\sigma$, with additional properties of monotonicity and sigmoidality, were algorithmically constructed in [5] and utilized in practical examples. It should be remarked that the existence of activation functions that ensure universal approximation for single and two hidden layer neural networks with a fixed number of hidden units was first established in [17].
If in Theorem 2.1, we take $\mathcal{A}(X)=\mathcal{K}(X)$ and any superactivation function $\sigma$, then the number of terms $r$ will be $12m+$ 1. To see this, it is sufficient to repeat the proof, noting that $n=2m+1$ and $k=1$. This observation leads to the following theorem.
Theorem 2.4. For $p=1,2,...,n$, let $X_{p}$ be a compact metric space of finite dimension $d_{p}$ and let $m=\sum_{p=1}^{n}d_{p}.$ There exist universal continuous functions $\psi_{pq}:X_{p}\rightarrow[0,1],$ $p=1,...,n,$ $q=1,...,2m+1,$ and an infinitely differentiable function $\sigma:\mathbb{R}\rightarrow\mathbb{R}$ such that for every continuous function $g$ defined on $X=\Pi_{p=1}^{n}X_{p}$ and any $\varepsilon>0$ there exist $w_{q},\theta_{q},\in\mathbb{R}$, $q=1,...,2m+1$, such that
$\left|g(x_{1},...,x_{n})-\sum_{q=1}^{2m+1}\sigma(w_{q}\sum_{p=1}^{n}\psi_{pq}(x_{p})-\theta_{q})\right|<\varepsilon,$
for all $(x_{1},...,x_{n})\in X$.
Note that in Theorem 2.4 the outer function $\sigma$ does not depend on $g$. The only parameters that depend on $g$ are the numbers $\theta_{q}$. The numbers $w_{q}$ can be taken to be equal and fixed once and for all. This is evident from the construction of $\sigma$ above (see (2.6), where $w$ is fixed for all $u$). For example, if we set $\alpha=b-a$, where $[a,b]$ is a closed interval containing all the sets $\Psi_{q}(X),$ where $\Psi_{q}(x_{1},...,x_{n})=\sum_{p=1}^{n}\psi_{pq}(x_{p})$, $q=1,...,2m+1$, then $w_{q}$ can all be taken to be equal to $1$.

### v2

@@
 Mathematics Subject Classifications: 41A30, 41A65, 68T05
-Keywords: feedforward neural network, universal approximation theorem, density, topological vector space, Tauber-Wiener function, Kolmogorov’s superposition theorem
+Keywords: feedforward neural network, universal approximation theorem, density, topological space, locally convex space, Tauber-Wiener function, Kolmogorov’s superposition theorem
 1. Introduction
@@
 The most basic form of an MLP features just one hidden layer. In this setup, each output neuron calculates a function expressed as
-$\sum_{i=1}^{r}c_{i}\sigma(\mathbf{w}^{i}\cdot\mathbf{x}-\theta_{i}),$ (1.1)1.1( 1.1 )
+$\sum_{i=1}^{r}c_{i}\sigma(\mathbf{w}^{i}\cdot\mathbf{x}-\theta_{i}),$ (1.1)
 where $\mathbf{x}=(x_{1},...,x_{d})$ represents the input vector, $r$ is the number of neurons in the hidden layer, $\mathbf{w}^{i}$ are weight vectors in $\mathbb{R}^{d}$, $\theta_{i}$ are thresholds, $c_{i}$ are coefficients, and $\sigma$ is the activation function, a real univariate function.
@@
 This principle states that a neural network with a single hidden layer can approximate any continuous function over a compact domain to any desired level of precision.
-Specifically, the set $span\{\sigma(\mathbf{w\cdot x}-\theta):\theta\in\mathbb{R},\mathbf{w}\in\mathbb{R}^{d}\}$, which comprises functions defined in the format of equation (1.1), is dense in $C(K)$ for every compact set $K\subset\mathbb{R}^{d}$.
+Specifically, the set of all functions defined in the format of equation (1.1), is dense in $C(K)$ for every compact set $K\subset\mathbb{R}^{d}$.
 Here $C(K)$ represents the space of real-valued continuous functions on $K$.
 This important result in neural network theory is known as the universal approximation theorem (UAT).
-Extensive research has investigated the UAT across various activation functions $\sigma$, examining how different choices influence the approximation capabilities of neural networks.
-The most general result in this area was obtained by Leshno, Lin, Pinkus and Schocken [14].
-They proved that a continuous activation function $\sigma$ possesses the UAP if and only if it is not a polynomial.
-This result demonstrates the effectiveness of the single hidden layer perceptron model across a wide range of activation functions $\sigma$.
-It should be noted that, the universal approximation theorem in [14] was shown to apply to a broader class of activation functions beyond just continuous ones, including activation functions that may have discontinuities on sets of Lebesgue measure zero.
-However, this paper will specifically concentrate on continuous activation functions.
-For a thorough, step-by-step proof of this theorem, refer to [19, 20].
-In the past, it was commonly accepted and highlighted in numerous studies that attaining the universal approximation property necessitate large networks with a substantial number of hidden neurons (see, e.g., [4, Chapter 6.4.1]).
-In the above-mentioned earlier works, the number of hidden neurons was regarded as unbounded.
-However, more recent research [5, 6, 7] has demonstrated that neural networks using certain non-explicit but practically computable activation functions can approximate any continuous function over any compact set to any desired level of accuracy, even with a minimal and fixed number of hidden neurons.
 Note that the inner product $\mathbf{w}^{i}\cdot\mathbf{x}$ in (1.1) represents a linear continuous functional on $\mathbb{R}^{d}$.
-Conversely, by Riesz representation theorem, every linear functional on $\mathbb{R}^{d}$ is of the form $\mathbf{w\cdot x}$, where $\mathbf{w}\in\mathbb{R}^{d}$ and $\mathbf{x}=(x_{1},...,x_{d})$ is the variable (see [21, Theorem 13.32]).
+Conversely, by Riesz representation theorem, every linear functional on $\mathbb{R}^{d}$ is of the form $\mathbf{w\cdot x}$, where $\mathbf{w}\in\mathbb{R}^{d}$ and $\mathbf{x}=(x_{1},...,x_{d})$ is the variable (see [30, Theorem 13.32]).
 Linear continuous functionals constitute a significant subclass in $C(\mathbb{R}^{d})$, denoted here by $\mathcal{L}(\mathbb{R}^{d})$.
@@
   • Input Layer: This layer consists of an element $x\in X$, where $X$ is an arbitrary topological space.
-[…]
+  • Hidden layer: Each neuron in the hidden layer takes the input $x$ from the input layer and applies a function $f\in\mathcal{A}(X)$ to $x$.
+This value is then multiplied by a weight $w$.
+A shift $\theta$ and then a fixed activation function $\sigma:$ $\mathbb$$\rightarrow R$} are applied to $f(x)$.
+The resulting value $\sigma(wf(x)-\theta)$ represents the output signal of the neuron.
   • Output layer: Each neuron in this layer receives weighted signals from each neuron in the hidden layer, sums them up, and produces the final output value.
-This architecture significantly extends the traditional feedforward neural networks.
-When $X=\mathbb{R}^{d}$ and $\mathcal{A}(X)=\mathcal{L}(\mathbb{R}^{d})$, the input $x$ represents a $d$-dimensional vector.
-In this very special case, the layer contains $d$ traditional neurons, each receiving an input signal $x_{1},x_{2},...,x_{d}\in\mathbb{R}$, respectively.
-Note that in the aforementioned architecture, the element $x\in X$ carries all the information of the input layer.
-This structure enables the network to accommodate a wide diversity of input types.
+This architecture represents a substantial generalization of classical feedforward neural networks.
+Traditional networks are recovered only in the specific case where $X=\mathbb{R}^{d}$ and $\mathcal{A}(X)=\mathcal{L}(\mathbb{R}^{d})$, yielding a structure with $d$ input neurons, each processing a real-valued component $x_{1},x_{2},...,x_{d}\in\mathbb{R}$.
+In the more general setting, however, the input $x\in X$ carries all information processed by the input layer as a whole.
+This generalized architecture allows the network to handle a wide variety of input types.
 In general, a single hidden layer TFNN computes a function of the form
-$\sum_{i=1}^{r}c_{i}\sigma(w_{i}f_{i}(x)-\theta_{i}),$ (1.2)1.2( 1.2 )
-[…]
-The aim of this paper is to show that for a broad a class activation functions $\sigma$, neural networks of the form presented in (1.2) can approximate any continuous function on a compact subset $K\subset X$ with arbitrary precision.
+$\sum_{i=1}^{r}c_{i}\sigma(w_{i}f_{i}(x)-\theta_{i}),$ (1.2)
+where $x\in X$ is the input, $f_{i}\in\mathcal{A}(X),$ $c_{i},w_{i},\theta_{i}\in\mathbb{R}$ are the parameters of the network, and $\sigma:$ $\mathbb$$\rightarrow R$} is a fixed activation function.
+The aim of this paper is to show that for a broad a class of activation functions $\sigma$, neural networks of the form presented in (1.2) can approximate any continuous function on a compact subset $K\subset X$ with arbitrary precision.
 In other words, the set
@@
 As an application of this result, we will derive an approximative version of the Kolmogorov superposition theorem (KST) for compact metric spaces, where outer (non-fixed and generally nonsmooth) functions are substituted with a fixed ultimately smooth function.
+Related literature: Extensive research has investigated the UAT across various activation functions $\sigma$, examining how different choices influence the approximation capabilities of neural networks.
+The most general result in this area was obtained by Leshno, Lin, Pinkus and Schocken [23].
+They proved that a continuous activation function $\sigma$ possesses the UAP if and only if it is not a polynomial.
+This result demonstrates the effectiveness of the single hidden layer perceptron model across a wide range of activation functions $\sigma$.
+It should be noted that, the universal approximation theorem in [23] was shown to apply to a broader class of activation functions beyond just continuous ones, including activation functions that may have discontinuities on sets of Lebesgue measure zero.
+However, this paper will specifically concentrate on continuous activation functions.
+For a thorough, step-by-step proof of this theorem, refer to [28, 29].
+In the past, it was commonly accepted and highlighted in numerous studies that attaining the universal approximation property necessitate large networks with a substantial number of hidden neurons (see, e.g., [10, Chapter 6.4.1]).
+In the above-mentioned earlier works, the number of hidden neurons was regarded as unbounded.
+However, more recent research [11, 12, 14] has demonstrated that neural networks using certain non-explicit but practically computable activation functions can approximate any continuous function over any compact set to any desired level of accuracy, even with a minimal and fixed number of hidden neurons.
 It should be noted that the UAP of neural networks operating between Banach spaces has been explored in various studies.
 For example, in [33], the fundamentality of ridge functions was established in a Banach space and subsequently applied to shallow networks with a sigmoidal activation function (see also [24]).
+Recall that a subset $S$ in a topological vector space $E$ is said to be “fundamental” if the linear span of $S$ is dense in $E$.
 In [5], the authors showed that any continuous nonlinear function mapping a compact set $V$ in a Banach space of continuous functions $C(K_{1})$ into $C(K_{2})$ can be approximated arbitrarily well by shallow feedforward neural networks.
@@
 In [7], the scope of this architecture was extended by proving several universal approximation theorems for quasi-Polish input and output spaces.
+The UAT for so-called functional input neural networks, defined on weighted spaces with output values in a Banach space, was proved in [6].
+Modifications of traditional neural networks and neural network architectures for mappings between non-Euclidean spaces, different from those considered here, have been discussed in several other papers.
+For example, in [19], the authors addressed the question: Which modifications to the input and output layers of a neural network preserve its universal approximation capabilities?
+They obtained sufficient conditions on topological spaces $X$ and $Y$, along with a pair of feature maps $\phi:X\rightarrow\mathbb{R}^{n}$ and $\rho:\mathbb{R}^{m}\rightarrow Y$, ensuring that if a family of functions $\mathcal{F}$ is dense in $C(\mathbb{R}^{n},\mathbb{R}^{m})$, then the set $\left\{\phi\circ f\circ\rho:f\in\mathcal{F}\right\}$ is dense in $C(X,Y)$, where density is understood with respect to the topology of uniform convergence on compact sets.
+The set $\mathcal{F}$ represents an arbitrary expressive neural network architecture.
+For example, $\mathcal{F}$ can be taken as the set of single hidden layer networks with a continuous non-polynomial activation function.
+As a continuation of [19] and motivated by developments in the mathematics of deep learning, the authors of [20] constructed universal function approximators for continuous maps between arbitrary Polish metric spaces $X$ and $Y$, using elementary functions between Euclidean spaces as building blocks.
+In [1], a new geometric deep learning model called geometric transformers was proposed.
+The authors demonstrated that these models can approximate Hölder continuous functions $f:K\rightarrow Y$, where $K$ is a compact subset of $\mathbb{R}^{d}$ and $Y$ is a suitable metric space.
 In [21], universal approximation theorems were obtained for neural operators (NOs) and mixtures of neural operators (MoNOs) acting between Sobolev spaces.
@@
 We hope that the results of this paper will stimulate further exploration of these neural networks, particularly with outputs from these and other general spaces.
+Comparison with other results: To the best of our knowledge, this is the first paper to explore the UAT for neural networks defined on a general topological space.
+No additional structures, such as linearity, metricity, completeness, or others, are required.
+The only restriction of our main result is that the output space must be the usual real line.
+The techniques utilized in this paper are not yet sufficient to generalize the results to neural networks with outputs in infinite-dimensional spaces.
+However, we hope this research will stimulate future work in this direction.
+As indicated in the literature review above, all previous generalizations of UAT involve general input-output correspondences where the input is a topological space, always with additional structure, and the output is the familiar $\mathbb{R}^{n}$, or the input space is the same as before, and the output is also a topological space with additional structure.
+In most cases, this additional structure consists of norms and inner products, which often make the underlying spaces Banach or Hilbert spaces.
+Recent studies have also involved Fréchet spaces (complete and metrizable locally convex topological vector spaces) and even quasi-Polish spaces (general topological spaces that admit the existence of a countable family of real-valued continuous functions that separate points).
+Clearly, the problem of function approximation by neural networks with outputs in $\mathbb{R}^{n}$ can be reduced to the case where the output lies in $\mathbb{R}$.
+This is because any $\mathbb{R}^{n}$-valued function can be treated as a vector of $n$ scalar-valued functions, each mapping into $\mathbb{R}$.
+Therefore, it suffices to construct a neural network that approximates each scalar component individually, using a single neuron in the output layer for each component.
+Taking this into account, in the case of finite-dimensional outputs, our results can be compared with earlier results that involve shallow networks as follows.
+Our main result employs Tauber-Wiener activation functions, the broadest class of activation functions for which UAT holds.
+These functions automatically (by definition) guarantee approximation with arbitrary accuracy in the univariate setting.
+Without this basic property, no UAT can be proven in multivariate or more abstract settings.
+Some classical and recent papers devoted to UAT in general spaces involve sigmoidal functions (see [24, 33]), weak sigmoidal functions — those that tend to 0 as $x\rightarrow$ -$\infty$, but not necessarily to $1$ as $x\rightarrow$ $\infty$ (see [35]) — discriminatory functions (see [4]), or activation functions with a separating property (see [7]), which is the infinite-dimensional counterpart to the well-known sigmoidal property for real univariate functions.
+The work [5] considers Tauber-Wiener functions, but the underlying input space is Banach, which is far from a general topological space.
+Note that the mildest known condition imposed on a topological space is “quasi-Polish” (see [7]).
+In summary, we claim that if the output space is $\mathbb{R}^{n}$, then the UAT proven in this paper is more general with respect to both the input space and/or activation functions compared to classical and the recent corresponding results.
+We also apply our main result to obtain an approximate version of Kolmogorov’s superposition theorem for compact metric spaces.
+Note that although much research has focused on the role of the classical Kolmogorov theorem in conventional neural networks, and many studies have been devoted to the generalization of neural networks to infinite-dimensional spaces, the extension of Kolmogorov’s theorem (due to Ostrand) has not yet been considered.
+In Ostrand’s theorem, the outer functions depend on the target function and may be poorly behaved.
+Our second result shows that these outer functions can be replaced with a single, ultimately smooth, universal function (not depending on the target function) to approximate every continuous function on a compact metric space with arbitrary precision.
 2. Main results
@@
 We say that a subclass $\mathcal{A}(X)\subset C(X)$ holds the $D$-property if the set
-$S=span\left\{u\circ v:u\in C(\mathbb{R}),v\in\mathcal{A}(X)\right\}$ (2.1)2.1( 2.1 )
+$S=span\left\{u\circ v:u\in C(\mathbb{R}),v\in\mathcal{A}(X)\right\}$ (2.1)
 is dense in $C(X)$.
@@
 Assume $X$ is a topological space, $\mathcal{A}(X)$ is a subclass of $C(X)$ with the $D$-property and $\sigma:\mathbb{R}\rightarrow\mathbb{R}$ is a TW function.
-Then for any $\varepsilon>0$, any compact set $K\subset X$ and any function $g\in C(K)$ there exist $r\in\mathbb{N}$, $f_{i}\in\mathcal{A}(X)$, $c_{i},w_{i},\theta_{i},\in\mathbb{R}$, $i=1,...,r$, such that
+Then for any $\varepsilon>0$, any compact set $K\subset X$ and any function $g\in C(X)$ there exist $r\in\mathbb{N}$, $f_{i}\in\mathcal{A}(X)$, $c_{i},w_{i},\theta_{i},\in\mathbb{R}$, i=1,,ri=1,...,r, such that
 $\max_{x\in K}\left|g(x)-\sum_{i=1}^{r}c_{i}\sigma(w_{i}f_{i}(x)-\theta_{i})\right|<\varepsilon.$
@@
 Take any $\varepsilon>0$, any compact set $K\subset X$ and any function $g\in C(K)$.
-Since $\mathcal{A}(X)$ has the $D$-property, there exist finitely many functions $u_{i}\in C(\mathbb{R})$ and $v_{i}\in C(X)$ such that
-$\left|g(x)-\sum_{i=1}^{n}u_{i}(v_{i}(x))\right|<\varepsilon/2,$ (2.2)2.2( 2.2 )
+Since $\mathcal{A}(X)$ has the $D$-property, there exist finitely many functions $u_{i}\in C(\mathbb{R})$ and $v_{i}\in\mathcal{A}(X)$ such that
+$\left|g(x)-\sum_{i=1}^{n}u_{i}(v_{i}(x))\right|<\varepsilon/2,$ (2.2)
 for all $x\in K$.
@@
 for all $t\in V$. Therefore,
-$\left|u_{i}(v_{i}(x))-\sum_{j=1}^{k_{i}}c_{ij}\sigma(w_{ij}v_{i}(x)-\theta_{ij})\right|<\varepsilon/2n$ (2.3)2.3( 2.3 )
-for each $i=1,...,n$, and all $x\in K$. It follows from (2.2) and (2.3) that
+$\left|u_{i}(v_{i}(x))-\sum_{j=1}^{k_{i}}c_{ij}\sigma(w_{ij}v_{i}(x)-\theta_{ij})\right|<\varepsilon/2n$ (2.3)
+for each i=1,,ni=1,...,n, and all $x\in K$. It follows from (2.2) and (2.3) that
 $\left|g(x)-\sum_{i=1}^{n}\sum_{j=1}^{k_{i}}c_{ij}\sigma(w_{ij}v_{i}(x)-\theta_{ij})\right|<\varepsilon$
@@
 Theorem 2.1 generalizes existing universal approximation theorems for traditional feedforward neural networks.
-This is because in traditional networks, the space of linear continuous functionals on $\mathbb{R}^{d}$ serves as the basic family $\mathcal{A}(X)$, which clearly satisfies the $D$-property.
+In such networks, the space of linear continuous functionals on $\mathbb{R}^{d}$ serves as the basic family $\mathcal{A}(X)$, which clearly satisfies the $D$-property.
+Indeed, linear continuous functionals on $\mathbb{R}^{d}$ are of the form $\mathbf{a}\cdot\mathbf{x}$, where $\mathbf{a}\in\mathbb{R}^{d}$, $\mathbf{x}\in\mathbb{R}^{d}$, and $\cdot$ denotes the scalar product.
+Since the linear span of ridge functions $g(\mathbf{a}\cdot\mathbf{x)}$ is dense in $C(\mathbb{R}^{d})$ in the topology of uniform convergence on compact sets (see, e.g., [14, p.
+12]), $D$-property is satisfied in this setting.
 Note that, in particular, $X$ may be a topological vector space.
@@
 Assume $X$ is a locally convex topological vector space (in particular, a normed space) and $\sigma$ is a continuous univariate function that is not a polynomial.
-Then for any $\varepsilon>0$, any compact set $K\subset X$ and any function $g\in C(K)$ there exist $r\in\mathbb{N}$, $f_{i}\in X^{\ast}$, $c_{i},\theta_{i}\in\mathbb{R}$, $i=1,...,r$, such that
+Then for any $\varepsilon>0$, any compact set $K\subset X$ and any function $g\in C(K)$ there exist $r\in\mathbb{N}$, $f_{i}\in X^{\ast}$, $c_{i},\theta_{i}\in\mathbb{R}$, i=1,,ri=1,...,r, such that
 $\max_{x\in K}\left|g(x)-\sum_{i=1}^{r}c_{i}\sigma(f_{i}(x)-\theta_{i})\right|<\varepsilon.$
@@
 Now since $X$ is locally convex, the Hahn-Banach extension theorem holds.
-It is a consequence of this theorem that for any distinct points $x_{1},x_{2}\in X$ there exists a functional $r\in X^{\ast}$ such that $r(x_{1})\neq r(x_{2})$ (see [22, Theorem 3.6]).
+It is a consequence of this theorem that for any distinct points $x_{1},x_{2}\in X$ there exists a functional $r\in X^{\ast}$ such that $r(x_{1})\neq r(x_{2})$ (see [31, Theorem 3.6]).
 Hence, the algebra $S$ separates points in $X$.
@@
 Let us now we apply Theorem 2.1 to derive an approximative version of the renowned Kolmogorov superposition theorem (KST) for compact metric spaces.
-KST [10] states that for the unit cube $\mathbb{I}^{d},~{}\mathbb{I}=[0,1],~{}d\geq 2,$ there exist $12d+$ 1 functions $\{s_{q}\}_{q=1}^{2d+1}\subset C(\mathbb{I}^{d})$ of the form
+KST [17] states that for the unit cube $\mathbb{I}^{d},~\mathbb{I}=[0,1],~d\geq 2,$ there exist $2d+1$ functions $\{s_{q}\}_{q=1}^{2d+1}\subset C(\mathbb{I}^{d})$ of the form
 $s_{q}(x_{1},...,x_{d})=\sum_{p=1}^{d}\varphi_{pq}(x_{p}),~\varphi_{pq}\in C(\mathbb{I}),~p=1,...,d,~q=1,...,2d+1,$
@@
 This surprising and deep result, which solved (negatively) Hilbert’s 13-th problem, has been improved and generalized in several directions.
-For detailed information about KST, including its refinements, various variants, and generalizations, see the monographs [9, Chapter 1] and [7, Chapter 4].
+For detailed information about KST, including its refinements, various variants, and generalizations, see the monographs [16, Chapter 1] and [14, Chapter 4].
+It is worth noting that KST initiated a new line of research in approximation theory — namely, the approximation of multivariate functions by various combinations of functions of fewer variables (see, e.g., [2, 3, 8, 9, 13, 34]).
 The relevance of KST to neural networks, along with its theoretical and computational aspects, has been extensively discussed in the neural network literature (see, e.g., [15] and references therein).
@@
 (Ostrand [27]).
-For $p=1,2,...,n$, let $X_{p}$ be a compact metric space of finite dimension $d_{p}$ and let $m=\sum_{p=1}^{n}d_{p}.$ There exist universal continuous functions $\psi_{pq}:X_{p}\rightarrow[0,1],$ $p=1,...,n,$ $q=1,...,2m+1,$ such that every continuous function $g$ defined on $\Pi_{p=1}^{n}X_{p}$ is representable in the form
-$g(x_{1},...,x_{n})=\sum_{q=1}^{2m+1}h_{q}(\sum_{p=1}^{n}\psi_{pq}(x_{p})),$
+For p=1,2,,np=1,2,...,n, let $X_{p}$ be a compact metric space of finite dimension $d_{p}$ and let $m=\sum_{p=1}^{n}d_{p}.$ There exist universal continuous functions $\psi_{pq}:X_{p}\rightarrow[0,1],$ p=1,,n,p=1,...,n, q=1,,2m+1,q=1,...,2m+1, such that every continuous function $g$ defined on $\Pi_{p=1}^{n}X_{p}$ is representable in the form
+$g(x_{1},...,x_{n})=\sum_{q=1}^{2m+1}h_{q}\left(\sum_{p=1}^{n}\psi_{pq}(x_{p})\right),$
 where $h_{q}$ are continuous functions depending on $g$.
-It follows from this theorem that for the metric space $X=\Pi_{p=1}^{n}X_{p}$ the family of $12m+$ 1 functions
-$\mathcal{K}(X)=\left\{\sum_{p=1}^{n}\psi_{pq}(x_{p}):q=1,...,2m+1\right\}$
-satisfies the $D$-property in $C(X)$.
-Let now $\sigma$ be a specific infinitely differentiable TW function with the property that, for any interval $[a,b]$, the set $\Lambda=\{\sigma(wx-\theta):w,\theta\in\mathbb{R}\}$ is dense in $C[a,b]$.
-Note that this is not the linear span of the functions $\sigma(wx-\theta)$, but rather a very narrow subclass of it.
-Such functions $\sigma$ do indeed exist.
-To show this, let $\alpha$ be any positive real number.
-Divide the interval $[\alpha,+\infty)$ into the segments $[\alpha,2\alpha],$ $[2\alpha,3\alpha],...$.
+Although substantial research has been conducted on generalizing neural networks to abstract spaces, Ostrand’s theorem has not yet been considered in this context.
+Like Kolmogorov’s original superposition theorem, Ostrand’s result can also be interpreted as a special case of a two-hidden-layer feedforward neural network model, where the inputs lie in a general compact metric space.
+We refer the reader to [15] for such a model when the input space is $\mathbb{R}^{n}$.
+From an application perspective, however, Theorem 2.3 is not suitable, as the outer functions $h_{q}$ are not fixed and depend on the target function $g$.
+In standard feedforward neural network theory, activation functions are typically fixed and do not depend on the functions being approximated.
+Our next theorem demonstrates that Ostrand’s result can be adapted to the neural network framework by replacing the outer functions $g_{q}$, which may be highly non-smoth, with a single infinitely differentiable universal activation function $\sigma$.
+This yields a model capable of approximating any continuous function defined on a compact metric space.
+The proof relies on Theorem 2.1 and the construction of specific activation functions, which we refer to as superactivation functions.
+A function $\sigma\in C(\mathbb{R})$ with the property that, for any interval $a,b]\subset\mathbb{R}$, the set $\Lambda=\{\sigma(wx-\theta):w,\theta\in\mathbb{R}\}$ is dense in $C[a,b]$ is called a superactivation function.
+Note that here $\Lambda$ is not the linear span of the functions $\sigma(wx-\theta)$, but rather a very narrow subclass of it.
+Superactivation functions indeed exist, which shows the following proposition.
+Proposition 2.1. There exist an infinitely differentiable superactivation function $\sigma$. Moreover, such a function can be constructed algorithmically.
+Proof.
+Let $\alpha$ be any positive real number.
+Divide the interval $\alpha,+\infty)$ into the segments $\alpha,2\alpha], [$$2\alpha,3\alpha],$....
 Let $\{p_{n}(t)\}_{n=1}^{\infty}$ be the sequence of polynomials with rational coefficients defined on $[0,1].$ We construct $\sigma$ in two stages.
-In the first stage, we define $\sigma$ on the closed intervals $[(2m-1)\alpha,2m\alpha],$ $m=1,2,...$ as the function
+In the first stage, we define $\sigma$ on the closed intervals $(2m-1)\alpha,2m\alpha],$ $m=1,2,$...
+as the function
 $\sigma(t)=p_{m}(\frac{t}{\alpha}-2m+1),\text{ }t\in[(2m-1)\alpha,2m\alpha],$
 or equivalently,
-$\sigma(\alpha t+(2m-1)\alpha)=p_{m}(t),\text{ }t\in[0,1].$ (2.4)2.4( 2.4 )
-In the second stage, we extend $\sigma$ to the intervals $(2m\alpha,(2m+1)\alpha),$ $m=1,2,...,$ and $(-\infty,\alpha)$, maintaining the $C^{\infty}$ property.
+$\sigma(\alpha t+(2m-1)\alpha)=p_{m}(t),\text{ }t\in[0,1].$ (2.4)
+In the second stage, we extend $\sigma$ to the intervals $(2m\alpha,(2m+1)\alpha),$ m=1,2,,m=1,2,..., and $(-\infty,\alpha)$, maintaining the $C^{\infty}$ property.
+To do this we use the smooth transition function
+$\beta(t)=\left\{\begin{array}[]{c}e^{-\frac{1}{t}}\text{ if }t>0,\\ 0\text{ if }t\leq 0,\end{array}\right$.
+Then the function
+$\omega(t):=\frac{\beta(\frac{t-2m\alpha}{\alpha})}{\beta(\frac{t-2m\alpha}{\alpha})+\beta(1-\frac{t-2m\alpha}{\alpha})}$
+has the following properties
+1) $\omega(t)=0$ for $t\leq 2m\alpha,$
+2) $\omega(t)=1$ for $t\geq(2m+1)\alpha,$
+3) $\omega(t)\in(0,1)$ for $t\in(2m\alpha,(2m+1)\alpha),$
+4) $\omega\in C^{\infty}(\mathbb{R}).$
+Now the function
+$\tau(t):=(1-\omega(t))\sigma(2m\alpha)+\omega(t)\sigma((2m+1)\alpha)$
+smoothly transitions from $\sigma(2m\alpha)$ to $\sigma((2m+1)\alpha)$ on the interval $(2m\alpha,(2m+1)\alpha)$.
+By the similar way, we can smoothly extend $\sigma$ to the half line $(-\infty,\alpha)$.
 For any univariate function $h\in C[0,1]$ and any $\varepsilon>0$ there exists a polynomial $p(t)$ with rational coefficients such that
@@
 for all $t\in[0,1].$ This together with (2.4) mean that
-$\left|h(t)-\sigma(\alpha t-s)\right|<\varepsilon,$ (2.5)2.5( 2.5 )
+$\left|h(t)-\sigma(\alpha t-s)\right|<\varepsilon,$ (2.5)
 for some $s\in\mathbb{R}$ and all $t\in[0,1].$
@@
 Indeed, let $u\in C[a,b]$, $\sigma$ be constructed as above and $\varepsilon$ be an arbitrarily small positive number.
-The transformed function $h(t)=u(a+(b-a)t)$ is well defined on $[0,1]$ and we can apply the inequality (2.5).
+The transformed function h(t)=u(a+(ba)t)h(t)=u(a+(b-a)t) is well defined on $[0,1]$ and we can apply inequality (2.5).
 Now using the inverse transformation $t=\frac{x-a}{b-a}$, we can write that
-$\left|u(x)-\sigma(wx-\theta)\right|<\varepsilon,$ (2.6)2.6( 2.6 )
+$\left|u(x)-\sigma(wx-\theta)\right|<\varepsilon,$ (2.6)
 for all $x\in[a,b]$, where $w=\frac{\alpha}{b-a}$ and $\theta=\frac{\alpha a}{b-a}+s$.
-We define activation functions $\sigma$ as superactivation functions if they satisfy (2.6) for any $u\in C[a,b]$, $\varepsilon>0$, and some $w,\theta\in\mathbb{R}$.
-These functions demonstrate that shallow networks can approximate univariate continuous functions with the minimal number of hidden neurons; in fact, a single hidden neuron is sufficient.
+Inequality (2.6) proves the proposition.
+Superactivation functions demonstrate that shallow networks can approximate univariate continuous functions with the minimal number of hidden neurons; in fact, a single hidden neuron is sufficient.
 Similar activation functions $\sigma$, with additional properties of monotonicity and sigmoidality, were algorithmically constructed in [11] and utilized in practical examples.
 It should be remarked that the existence of activation functions that ensure universal approximation for single and two hidden layer neural networks with a fixed number of hidden units was first established in [26].
-If in Theorem 2.1, we take $\mathcal{A}(X)=\mathcal{K}(X)$ and any superactivation function $\sigma$, then the number of terms $r$ will be $12m+$ 1.
-To see this, it is sufficient to repeat the proof, noting that $n=2m+1$ and $k=1$.
-This observation leads to the following theorem.
 Theorem 2.4.
-For $p=1,2,...,n$, let $X_{p}$ be a compact metric space of finite dimension $d_{p}$ and let $m=\sum_{p=1}^{n}d_{p}.$ There exist universal continuous functions $\psi_{pq}:X_{p}\rightarrow[0,1],$ $p=1,...,n,$ $q=1,...,2m+1,$ and an infinitely differentiable function $\sigma:\mathbb{R}\rightarrow\mathbb{R}$ such that for every continuous function $g$ defined on $X=\Pi_{p=1}^{n}X_{p}$ and any $\varepsilon>0$ there exist $w_{q},\theta_{q},\in\mathbb{R}$, $q=1,...,2m+1$, such that
-$\left|g(x_{1},...,x_{n})-\sum_{q=1}^{2m+1}\sigma(w_{q}\sum_{p=1}^{n}\psi_{pq}(x_{p})-\theta_{q})\right|<\varepsilon,$
+For p=1,2,,np=1,2,...,n, let $X_{p}$ be a compact metric space of finite dimension $d_{p}$ and let $m=\sum_{p=1}^{n}d_{p}.$ Assume $\sigma:\mathbb{R}\rightarrow\mathbb{R}$ is a superactivation function.
+Then there exist universal continuous functions $\psi_{pq}:X_{p}\rightarrow[0,1],$ p=1,,n,p=1,...,n, q=1,,2m+1,q=1,...,2m+1, such that for every continuous function $g$ defined on $X=\Pi_{p=1}^{n}X_{p}$ and any $\varepsilon>0$ there exist $w_{q},\theta_{q},\in\mathbb{R}$, q=1,,2m+1q=1,...,2m+1, satisfying the inequality
+$\left|g(x_{1},...,x_{n})-\sum_{q=1}^{2m+1}\sigma\left(w_{q}\sum_{p=1}^{n}\psi_{pq}(x_{p})-\theta_{q}\right)\right|<\varepsilon,$
 for all $(x_{1},...,x_{n})\in X$.
+The proof is based on the preceding material. It follows from Theorem 2.3 that for the metric space $X=\Pi_{p=1}^{n}X_{p}$, the family of $2m+1$ functions
+$\mathcal{K}(X)=\left\{\sum_{p=1}^{n}\psi_{pq}(x_{p}):q=1,...,2m+1\right\}$
+satisfies the $D$-property in $C(X)$.
+Now, if in Theorem 2.1 we take $\mathcal{A}(X)=\mathcal{K}(X)$ and choose any superactivation function $\sigma$, then the number of terms $r$ will be $2m+1$.
+To see this, it is sufficient to repeat the proof of Theorem 2.1, noting that $n=2m+1$, $k_{i}=1$ and all $c_{ij}=1$.
+This observation completes the proof of Theorem 2.4.
 Note that in Theorem 2.4 the outer function $\sigma$ does not depend on $g$.
+Moreover, by Proposition 2.1, $\sigma$ can be chosen to be infinitely differentiable.
 The only parameters that depend on $g$ are the numbers $\theta_{q}$.
@@
 This is evident from the construction of $\sigma$ above (see (2.6), where $w$ is fixed for all $u$).
-For example, if we set $\alpha=b-a$, where $[a,b]$ is a closed interval containing all the sets $\Psi_{q}(X),$ where $\Psi_{q}(x_{1},...,x_{n})=\sum_{p=1}^{n}\psi_{pq}(x_{p})$, $q=1,...,2m+1$, then $w_{q}$ can all be taken to be equal to $1$.
+For example, if we set $\alpha=b-a$, where $[a,b]$ is a closed interval containing all the sets $\Psi_{q}(X),$ where $\Psi_{q}(x_{1},...,x_{n})=\sum_{p=1}^{n}\psi_{pq}(x_{p})$, q=1,,2m+1q=1,...,2m+1, then $w_{q}$ can all be taken to be equal to $1$.
+Remark.
+The concept of topological space feedforward neural networks (TFNNs) and Theorem 2.1, our main result on the universal approximation property, are presented in the general framework of a topological space, without assuming the existence of a vector structure on the input set.
+This level of generality allows the theory to apply to a wide range of spaces lacking addition and scalar multiplication.
+In particular, if the input space $X$ is assumed to be a locally convex topological vector space, then Theorem 2.1 yields Theorem 2.2.
+Similarly, when $X$ is a product of compact metric spaces, the result leads to Theorem 2.4.
+Hence, these two cases arise as natural subclasses of the general framework considered here.
+Acknowledgements.
+The authors would like to thank the anonymous reviewers for their careful reading of the manuscript and their insightful comments and suggestions, which have helped to improve the clarity and quality of the paper.